### v1

# Chat2SVG: Vector Graphics Generation with Large Language Models and Image Diffusion Models

Abstract
Scalable Vector Graphics (SVG) has become the de facto standard for vector graphics in digital design, offering resolution independence and precise control over individual elements. Despite their advantages, creating high-quality SVG content remains challenging, as it demands technical expertise with professional editing software and a considerable time investment to craft complex shapes. Recent text-to-SVG generation methods aim to make vector graphics creation more accessible, but they still encounter limitations in shape regularity, generalization ability, and expressiveness. To address these challenges, we introduce Chat2SVG, a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. Our approach first uses an LLM to generate semantically meaningful SVG templates from basic geometric primitives. Guided by image diffusion models, a dual-stage optimization pipeline refines paths in latent space and adjusts point coordinates to enhance geometric complexity. Extensive experiments show that Chat2SVG outperforms existing methods in visual fidelity, path regularity, and semantic alignment. Additionally, our system enables intuitive editing through natural language instructions, making professional vector graphics creation accessible to all users. The project page is https://chat2svg.github.io.
1 Introduction
Scalable Vector Graphics (SVG), a vector image format based on geometric shapes, has become the standard for modern digital design, offering resolution independence and precise control over individual elements. However, creating high-quality SVG content remains challenging for non-expert users, as it requires both expertise with professional design software and considerable time investment to create complex shapes. To make vector graphics creation more accessible, recent research has focused on developing text-to-SVG systems that enable users to express their creative ideas through simple text prompts rather than complex manual editing.
Existing approaches have explored image-based methods for text-to-SVG generation, iteratively optimizing a large collection of shape elements (e.g., cubic Bézier curves, typically to ) by rendering them into images with differentiable rasterizers [20] and evaluating them using text-image similarity metrics like CLIP loss [28] and Score Distillation Sampling (SDS) loss [27]. While these image-based methods [9, 16, 43, 44, 50] can generate visually impressive SVG through the combination of numerous paths and powerful image models, they face a critical limitation in maintaining the regularity and semantics of shapes. Specifically, semantic components that should be represented as single elements often end up fragmented across multiple overlapping paths. Although these fragmented paths, when viewed collectively, can yield visually appealing outputs, they fundamentally conflict with professional design principles, where each semantic component is intentionally crafted as a single, regularized path.
Given SVG is defined using Extensible Markup Language (XML), another solution to text-to-SVG synthesis involves language-based methods, which treat SVG scripts as text input. Recent works [41, 35] have proposed specialized tokenization strategies to embed SVG scripts and trained Sequence-To-Sequence (seq2seq) models on domain-specific datasets (e.g., icons and fonts). While these approaches achieve good generation quality within their training domains, they suffer from limited generalization due to the absence of large-scale, general-purpose text-SVG training data. Meanwhile, some works [26, 31] show that while state-of-the-art Large Language Models (LLMs) can generate basic geometric shapes (e.g., circles and rectangles) with layouts matching text prompts, they struggle to produce the complex geometric details required for professional SVG applications. Consequently, the limited generalization and poor expressiveness of language-based methods hinder their adoption for text-to-SVG generation.
To address the aforementioned problems, we propose a novel hybrid framework that leverages the complementary strengths of image-based and language-based methods. Our approach first utilizes an LLM to synthesize SVG templates composed of basic primitives and then optimizes their geometric details guided by image diffusion models. This hybrid framework addresses the key limitations of both paradigms: The LLM-based template generation overcomes the domain-specific generalization constraints of traditional language models while ensuring shape regularity, as each element naturally corresponds to a single semantic component. The subsequent image-based optimization then enhances the expressiveness of these well-structured templates by capturing complex geometric details that LLMs alone struggle to produce. Specifically, to fully exploit LLMs’ ability to create reasonable and complex layouts that match text prompts, we design an SVG-oriented prompt with multiple stages, including prompt expansion, SVG script generation, and visual refinement. To refine the generated SVG templates, we implement a dual-stage optimization strategy: (1) We first use an image diffusion model to synthesize detailed image-level targets, then follow [50] to optimize path-level latent vectors with a pretrained SVG VAE, eliminating common issues such as self-intersections and jagged curves; (2) We further refine geometric details by directly optimizing the point coordinates to capture fine-grained visual elements. Comprehensive experiments show that Chat2SVG consistently outperforms existing methods regarding overall visual quality, individual path regularity, and semantic coherence. Furthermore, our system enables intuitive editing through iterative natural language instructions, making vector graphics creation more accessible to non-expert users. Our contributions are summarized as follows:
Hybrid SVG Generation Framework. We introduce a novel text-to-SVG generation paradigm that combines Large Language Models with image diffusion models to produce high-quality SVG outputs.
SVG-Oriented Prompt Design. We develop a specialized prompt system that directs LLMs to generate SVG templates using basic geometric primitives.
Dual-Stage Optimization. We implement a two-phase optimization process that preserves the semantic meaning of each shape while eliminating artifacts such as self-intersecting and jagged paths.
Iterative Editing. We enable iterative refinement of SVG through natural language instructions, making SVG creation more accessible.
2 Related Work
2.1 Large Language Models for Design
LLMs have emerged as powerful tools for graphics design tasks, demonstrating remarkable capabilities in understanding and reasoning about complex design specifications across diverse domains. Recent works have explored their applications in (1) visual layout and composition, including poster designs [34, 5, 47, 21], 3D scene arrangements [38, 14, 33, 51, 32, 2, 10], and placement of basic shapes for image generation [8, 46]; (2) content creation and manipulation for shapes [11], materials [15], and animation [22, 36]; and (3) design understanding [18, 42]. While these demonstrate LLMs’ versatility, their potential for vector graphics remains unexplored. Our work addresses this gap by showing that LLMs can effectively generate structured SVG templates, extending their capabilities to the domain of vector graphics creation.
2.2 Text-Guided Vector Graphics Generation
Text-to-SVG generation approaches can be roughly categorized into image-based and language-based methods. Image-based methods [9, 16, 43, 44, 50] start with a large collection of randomly initialized shapes, render vector graphics with a differentiable renderer [20], and iteratively optimize path parameters by minimizing text-image similarity losses (e.g., CLIP [28] loss and Score Distillation Sampling [27] loss). Despite their visually pleasant appearance, these methods often generate SVG results containing multiple fragmented paths that lack individual semantic correspondence, which is undesirable for real design scenarios. Language-based methods [41, 3, 35] design specialized tokenization strategies to encode vector graphics into discrete tokens, which are then concatenated with text tokens, and train seq2seq models like autoregressive transformers [37] on domain-specific datasets (e.g., icons and fonts). While these approaches are conceptually elegant, their generalization ability is limited by the scarcity of large-scale, general-purpose vector graphics datasets. Some works [26, 31] have explored LLMs’ vector graphics generation capabilities, showing that while LLMs can create reasonable layouts matching text prompts, they struggle to produce complex geometric shapes. Our Chat2SVG combines image-based and language-based methods to address the limitations of each paradigm, producing semantically meaningful SVG where each path corresponds to a distinct visual element and contains refined geometric details.
2.3 Vector Graphics Representation Learning
Vector graphics representation learning is essential for sketch-based retrieval, reconstruction, and generation tasks. A seminal study, SketchRNN [12], combines a Recurrent Neural Network (RNN) and a VAE to learn vector representations. Lopes et al. [24] trained a VAE to represent image-level font styles, then passed the latent embedding into a decoder for vector font generation. To increase reconstruction quality, later approaches like Sketchformer [30] and DeepSVG [4] adopted Transformer [37] architectures, followed by dual-modality methods [39, 23, 40] that leveraged both vector and image features. However, encoding the entire complex SVG into a single latent embedding often causes detail loss. To address this limitation, Zhang et al. [50] proposed path-level representational learning, creating a smooth latent space for individual paths. We leverage these path-level latent embeddings in our optimization to eliminate self-intersecting and jagged paths.
3 Method
Our Chat2SVG, as shown in Figure 2, begins with an SVG-oriented prompting approach (Section 3.1) that guides an LLM to create reasonable SVG templates. We then perform a dual-stage optimization (Section 3.2) to enhance the geometric details of SVG templates by optimizing SVG paths in both latent and control point space, guided by image diffusion models.
3.1 SVG Template Generation with LLMs
Prompt Expansion. Users often provide vague and brief prompts to the LLM when describing their desired graphics. Such unstructured input may lead to oversimplified SVG (regarding both the number of elements and structures) that fails to reflect the intended design. Consequently, we propose a three-layer prompt expansion strategy that systematically delineates the initial prompts: (1) Scene-level: Starting with an initial prompt, we instruct the LLM to analyze it holistically and identify essential objects that should appear. The LLM then expands the prompt by suggesting complementary objects to enhance scene completeness. (2) Object-level: For each object in the expanded prompt, we guide the LLM to systematically break it down into its components. For instance, when describing a lion, the LLM deconstructs it into distinct parts such as the body, head, mane, legs, eyes, ears, and tail. This decomposition ensures that no critical components are missing during SVG script generation. (3) Layout-level: After object decomposition, we direct the LLM to develop a comprehensive layout plan. This includes determining the position and size of each element on the canvas, selecting appropriate colors for visual harmony, and establishing clear spatial relationships to ensure a cohesive overall composition.
SVG Script Generation. After prompt expansion, we obtain a detailed scene description. We then convert this natural language specification into SVG scripts in XML format. Since the LLM has limitations in synthesizing geometrically complex paths, we constrain the shapes to a carefully selected set of basic primitives: rectangles, ellipses, lines, polylines, polygons, and short paths. We fix the canvas size to . Each geometric element is assigned a unique ID and includes semantic annotations in its comments.
Visual Rectification. Since prompt expansion is purely text-based, even when the generated SVG script accurately follows the detailed prompt, visual inconsistencies (e.g., misaligned components, disproportionate scaling, and incorrect path ordering) can emerge during rendering. For example, in Figure 2, the initial SVG template shows a lion with a misshapen mane and missing facial features. We thus adopt a visual rectification strategy in which we render the SVG and provide the rendered image back to the LLM (with vision capability) for inspection. The LLM analyzes the visual output, identifies any inconsistencies or oddities, and generates corrected SVG code. This visual refinement loop can be performed iteratively. In our experiments, we find that two iterations of refinement are typically sufficient to generate well-structured SVG templates with appropriate spatial layouts.
Furthermore, to enhance the quality of prompt expansion and SVG script generation, we provide curated in-context examples in the prompts. The complete set of prompts and examples is available in the supplementary material.
3.2 SVG Optimization Guided by Image Diffusion
Detail Enhancement. While the primitive shapes in the SVG template have accurate semantic meaning, they lack the sophisticated geometric details necessary for a professional design. To strengthen visual expressiveness, we first render the SVG template into a raster image. We then use an image editing method to generate a more detailed version that serves as our optimization target. Specifically, we employ SDEdit [25], which enhances the input image by first adding noise and then progressively denoising it with an image diffusion model to produce outputs with richer details, such as the contour of the lion’s mane illustrated in Figure 2. To maintain structural similarity between the enhanced output and the original image, we incorporate a ControlNet (tile version) [49] into our pipeline. By using Gaussian-blurred versions of the initial images as control signals, the ControlNet effectively maintains the overall composition throughout the enhancement process. Apart from geometric details, this image-to-image translation process also introduces new semantic parts such as the lion’s beard. To faithfully reproduce the target image, we propose using the Segment Anything Model (SAM) [17] to identify and add these new parts into the SVG template. In particular, we denote the rendered images of the initial SVG template and its diffusion-enhanced version as and , respectively. We apply SAM to both images to obtain their corresponding mask sets: for the template image and for the target image, where and represent the number of masks in each set. The major idea of path addition is that the diffusion process primarily introduces small decorative details (e.g., the beard and toes of the lion in Figure 2) while maintaining the key components (e.g., the tail and the body). Therefore, we implement a mask-filtering strategy to identify these new elements: For each mask in , we compare it against all masks in using Intersection over Union (IoU) loss [29]. If overlaps with a template mask and their IoU is below a threshold, we identify as a new decorative detail that should be added to the SVG. Conversely, if their IoU exceeds the threshold, we interpret as a shape variation of an existing component and exclude it from further processing. Finally, for all filtered masks, we approximate their boundaries with polygons and add them to the SVG script.
Notation. We define an SVG script as a collection of paths (i.e., shapes), . Each path consists of a sequence of commands, . A command is defined by its type and its associated control points , where M represents Move and C represents cubic Bézier curves. For consistency, we convert all other primitive shapes (e.g., rectangles and ellipses) in the SVG template into cubic Bézier curves.
Previous image-based methods [9, 16, 43, 44] optimize SVG shapes by directly manipulating the control points . Despite convenient, this approach often leads to artifacts such as self-intersecting curves and unnatural deformations. To overcome these limitations, Zhang et al. [50] introduced a path-level SVG VAE. This model’s latent space effectively captures common shape patterns and geometric constraints, allowing for the optimization of path latent vectors at a higher level to produce smooth outputs.
Latent Optimization. We leverage this pretrained SVG VAE to conduct the latent optimization. However, their SVG encoder expects a fixed number of commands ( cubic Bézier curves), while our SVG primitives contain varying, typically fewer commands. To resolve this mismatch, we develop a latent inversion process that converts our primitives into latent embeddings. We start with a randomly sampled latent vector and decode it into a command sequence. Then we evenly sample points along its contour. Similarly, we sample an equal number of points along the contour of our target primitive shape. Using these point sets, we compute the Earth Mover’s Distance (EMD) [1, 7, 48]:
where represents a bijective mapping between the point sets. By minimizing this EMD loss, we gradually optimize the latent embedding to match the target primitive.
After obtaining latent vectors for all paths, we jointly optimize these vectors along with visual attribute parameters: filling colors , stroke properties (color and width) , and transformation matrices . At each optimization iteration , we decode the latent vectors into command sequences and apply transformation matrices to the control points. We then combine color and stroke attributes together, and use a differentiable rasterizer [20] to render the complete SVG into an image . Our optimization objective consists of three loss terms: (1) An MSE loss between the rendered image and target image to ensure visual similarity; (2) A curvature loss that reduces sharp bends and fluctuations along the contour by computing the discrete second derivative:
(3) A path-level IoU loss that preserves the overall composition and prevents paths from drifting too far from their initial positions:
where and denote the binary masks of the -th path at the current iteration and its initial state, respectively.
The final loss function combines these three terms with empirically determined weights:
where and .
Point Optimization. While latent optimization effectively aligns SVG shapes with their target positions and contours, it has a notable limitation: the resulting shapes tend to be overly smooth, lacking the intricate details that are often essential in professional vector graphics. For example, in Figure 2, the lion’s mane after latent optimization does not capture the fine contour present in the target image. Therefore, we add a second optimization stage that directly refines the control points of the paths, similar to previous image-based works [9, 16, 43, 44]. To achieve finer granularity in shape control, we split each cubic Bézier curve at its midpoint, effectively doubling the number of control points in the SVG. This increased control point density allows for more precise shape adjustments. The point optimization stage employs a loss function that combines the MSE loss and the curvature loss:
The parameter decreases linearly from to throughout the optimization process.
3.3 Iterative Editing
During SVG script generation (Section 3.1), we instruct the LLM to annotate the semantic label of each path. This design enables users to refine the initial SVG template using natural language, as the LLM can precisely locate and modify specific elements based on semantic understanding.
Our system supports iterative refinement through multiple rounds of natural language instructions. However, across different iterations, when passing the SVG template image through SDEdit, the enhanced images may contain varying decorative details, potentially leading to inconsistencies in the final optimized SVG. To maintain consistency across editing iterations, we instruct the LLM to output both the modified SVG script and a list of specifically changed paths. This precise tracking enables selective optimization of modified shapes while preserving unaltered elements from previous iterations, maintaining visual coherence throughout the entire refinement process.
4 Experiments
Implementation Details. In our experiments, we use claude-3-5-sonnet as our backbone LLM model due to its leading generation capabilities. To evaluate our approach, we create a prompt set by having the LLM generate text prompts across five categories: animals, food, objects, scenes, and novel concept combinations. For each prompt, we generate one SVG template and perform rounds of visual rectification. This process repeats times, yielding candidate SVG templates per prompt. These templates are rendered as images, and we follow a standard practice to select the highest quality SVG using the ImageReward [45] metric. Detailed optimization configurations are provided in the supplementary materials.
Baselines. We compare our approach against five state-of-the-art image-based text-to-SVG generation methods: CLIPDraw [9], DiffSketcher [43], VectorFusion [16], SVGDreamer [44], and T2V-NPR [50]. Current language-based methods [41, 35] are confined to specific categories, with no support for general text-to-SVG generation, so we do not include them in our comparison. For a fair comparison, we set these baseline methods to use a similar number of shapes as our optimized SVG.
Evaluation Metrics. We evaluate the quality of our generated SVG across three dimensions: visual fidelity, path regularity, and semantic alignment.
Image-level Fidelity. We evaluate visual fidelity using a ground-truth dataset of well-designed vector graphics, specifically colored SVG files downloaded from SVGRepo111https://www.svgrepo.com. We compute the Fréchet Inception Distance (FID) [13] between the rendered images of our generated SVG and this professional design collection, using features extracted by the CLIP image encoder [28].
Path-level Regularity. We assess path quality using a transformer-based Auto-Encoder trained on the FIGR-8-SVG [6] dataset, following the DeepSVG [4] architecture. Unlike previous methods [50] that need a fixed number of commands, our model encodes each drawing command into a latent vector, accommodating varying command lengths. After training, we compute the mean of command latent embeddings for each path and calculate the FID between these and the ground truth paths from FIGR-8-SVG. This quantifies how well our paths align with professional SVG design patterns.
Text-level Alignment. We assess semantic alignment by computing the CLIP score [28] between the text prompt and the rendered SVG image.
4.1 Comparison with Existing Methods
Quantitative Comparison. Table 1 shows the evaluation metrics for all baseline methods. Our method achieves the best image FID score, indicating that our generated SVG closely aligns with professional design patterns. Regarding path regularity, our method yields paths that are closest to the professionally designed SVG, as evidenced by the lowest path FID. Furthermore, our Chat2SVG achieves the highest score in text-SVG alignment, validating the significance of LLM-generated SVG templates.
| Method |
| CLIPDraw [9] | |||||||||
| DiffSketcher [43] | |||||||||
| VectorFusion [16] | |||||||||
| SVGDreamer [44] | |||||||||
| T2V-NPR [50] | |||||||||
| Chat2SVG (ours) |
Qualitative Comparison. In Figure 3, we present a side-by-side comparison between our Chat2SVG and the baselines. Methods that optimize open-ended strokes based on CLIP or Diffusion models (i.e., CLIPDraw [9] and DiffSketcher [43]) produce results with messy visual appearances and poor text alignment. These methods sometimes fail to synthesize complete objects, such as the flamingo (first row) and the plane (third row) in Figure 3. In contrast, our method generates SVG results with clear layouts and greater fidelity to the prompts.
Methods that optimize closed shapes via score distillation of diffusion models (i.e., VectorFusion [16], SVGDreamer [44], and T2V-NPR [50]) can better align with the input text, but still produce fragmented paths with limited semantic meaning. For instance, in the plane example (third row of Figure 3), VectorFusion and SVGDreamer create planes and clouds composed of multiple jagged, irregular, and fragmented shapes, which only appear meaningful when viewed as a whole. This hinders convenient editing by graphic designers. T2V-NPR [50] addresses the issue of jagged paths using a path VAE (although it can be overly smooth at times) and reduces fragmented shapes by merging shapes with similar colors. However, its merging operation ignores the semantic meaning of shapes, resulting in the body of the plane being merged with the surrounding clouds. Additionally, the semantic meaning of the path is sometimes unclear, as exemplified by the cloud in the bottom-right corner, whose contour is actually approximated by the blue background instead of a true cloud shape. In contrast, our Chat2SVG adopts a unique approach by using LLMs to generate paths representing semantic components and performing dual-stage optimization based on image diffusion to enhance path expressiveness, thus ensuring both path regularity and visual fidelity of generated SVG.
User Study. To further compare Chat2SVG with baseline methods, we conducted a user study examining the same key aspects: (1) overall visual aesthetics, (2) path regularity, and (3) text-SVG alignment. We randomly sampled prompts and generated SVG outputs using our approach and five baseline methods. For each prompt, participants were shown all generated SVG results along with the corresponding text prompt and highlighted paths. They were then asked to select the highest quality result regarding the three evaluation criteria. We recruited participants ( male, female) through university mailing lists, with a mean age of years. Among them, participants reported prior experience in graphic design, providing a balanced mix of expert and novice perspectives. Analyzing the average selection ratio across all three metrics (Figure 4), we found that Chat2SVG-generated results were consistently preferred by participants over baseline methods.
4.2 Ablation Study
We conduct ablation experiments to evaluate the effectiveness of key components in our pipeline (Section 3.1). First, we remove the SVG template generation and use randomly initialized shapes to approximate target images. As shown in the second row of Table 2, both image FID and path FID values increase substantially. This performance degradation can be attributed to the substantial gap between randomly initialized shapes and target shapes, making optimization more challenging. Meanwhile, in the second column of Figure 5, the dog’s legs and ears lack semantic-clear components, highlighting the importance of SVG template generation. Second, we eliminate the dual-stage optimization. As shown in the third row of Table 2, this results in the lowest path FID, as primitives generated by LLMs naturally align with the path regularity presented in the ground-truth dataset. However, the visual outcome in Figure 5 lacks the necessary details, indicating that the overall visual quality suffers without dual-stage optimization. A more detailed ablation study is provided in the supplementary material.
| Setup |
| Full | |||||||||
| No SVG Template | |||||||||
| No SVG Optimization |
4.3 Iterative Editing Results
In this section, we present the iterative editing results. As demonstrated in Figure 6, users can progressively refine the SVG through multiple rounds of natural language instructions. During editing, we preserve the optimized shapes from previous iterations while only modifying the specified shapes. The results showcase how Chat2SVG can accurately interpret and execute user editing requests, maintaining both semantic coherence and visual quality throughout the modifications. Additional editing examples are provided in the supplementary material.
5 Conclusion
In this paper, we present Chat2SVG, a novel paradigm for text-to-SVG generation that combines LLMs and image diffusion models. Through our carefully designed SVG template generation and dual-stage SVG optimization pipeline, our method generates high-quality SVG outputs that exhibit strong visual fidelity, path regularity, and text alignment.
Despite our method’s superior results, there are several aspects that could be further improved: (1) LLMs may produce imprecise visual layouts, such as the burger in Figure 7. This can be alleviated by collecting an SVG dataset and extracting the corresponding layouts to fine-tune LLMs. (2) Our approach uses SAM to identify new parts. However, SAM may overlook important regions, like the pink frosting on the donut in Figure 7. Incorporating semantic-aware SAM [19] could potentially produce more accurate results. (3) The pretrained SVG VAE from [50] has a fixed number of commands. This makes it hard to represent intricate shapes. For example, the lion’s mane in Figure 7 lacks sufficient detail. Training a new SVG VAE with varying-length commands would help solve this problem. Our work can be extended to generate other types of vectorized data, such as icons and fonts. Additionally, through consistent editing of SVG across multiple frames, coupled with shape interpolation, our Chat2SVG has the potential to facilitate keyframe animation generation.
Supplementary Material
Appendix A Overview
This supplementary material provides additional implementation details and experimental results, including:
More details about the optimization procedure (Section B);
Additional ablation studies evaluating the effects of prompt design strategies, SAM-guided path additions, and various optimization stages (Section C);
Complete LLM prompts used in SVG generation and editing (Section D);
More generated and edited SVG examples (Section E) are available on the project website.
Appendix B Optimization Details
Our SVG optimization consists of three sequential stages: latent inversion, latent optimization, and point optimization. Each stage runs for iterations. During the latent optimization stage, we initialize the stroke color to black with a width of . We also set the gradient of the RGBA color’s alpha channel to zero to forbid transparency optimization. Upon completing latent optimization, we apply transformation matrices to the points, which allows us to bypass the need for optimizing transformation matrices in the subsequent point optimization stage. Optimizing an SVG containing approximately shapes requires around minutes when executed on a single NVIDIA RTX GPU, consuming roughly GB of GPU memory.
Appendix C Additional Ablation Study
SVG Template. We demonstrate the critical role of each component in SVG template generation, with results shown in Figure 8. (1) The prompt expansion phase is crucial for synthesizing complete SVG templates by enriching the initial brief description with essential details and components. Without it, key visual elements are missing, such as the dog’s chef hat and the police car’s flashing lights, leading to incomplete and semantically deficient outputs. This demonstrates how prompt expansion helps capture essential visual elements needed for a coherent design. (2) Without visual rectification, the generated SVG contains significant visual inconsistencies that impact both local and global coherence. At the local level, we observe misalignments between connected components (like the dog’s ear and head), while at the global level, certain elements become unrecognizable (such as the police car’s body) due to improper spatial relationships and proportions. These issues highlight how visual rectification serves as a crucial quality control step in ensuring the generated SVG maintains proper visual structure and semantic clarity.
Detail Enhancement. The SAM-guided path addition enhances the visual richness of the optimized SVG by incorporating decorative elements and fine details. Without this step, the generated SVG appears simpler. As shown in the fourth column of Figure 8, compared to the complete method in the first column, decorative elements such as the intricate patterns on the dog’s chef hat and the detailed windows of the police car are absent, resulting in reduced visual sophistication and artistic appeal.
SVG Optimization. Our experiments validate the importance of each optimization stage: (1) Without latent optimization, relying solely on point optimization produces shapes with undesirable artifacts like self-intersections in the dog’s hat and distorted corners in the car’s window; (2) Conversely, omitting point optimization yields excessively smooth contours in elements like the hat and window, resulting in a lack of geometric precision. We further analyze the impact of individual loss terms. The curvature loss plays a vital role in maintaining smooth and natural path contours - when excluded, the optimization generates shapes with noticeably irregular boundaries. Similarly, the path IoU loss is crucial for spatial consistency, anchoring elements in their designated positions. Without this constraint, components like the dog’s ear and window frame drift from their intended locations, compromising the overall compositional integrity.
Appendix D LLM Prompts
[…]
Appendix E More Results
To provide a more user-friendly visualization, we have put all generated SVG, including text-guided generation and editing results, on the project page.
| System Prompt |
| You are a vector graphics designer tasked with creating Scalable Vector Graphics (SVG) from a given text prompt. Your tasks: • Task 1: **Expand Text Prompt**. The provided prompt is simple, concise and abstract. The first task is imagining what will appear in the image, making it more detailed. • Task 2: **Write SVG Code**. Using the expanded prompt as a guide, translate it into SVG code. • Task 3: **Code Improvement**. Although the SVG code may align with the text prompt, the rendered image could reveal oddities from human perception. Adjust the SVG code to correct these visual oddities. Constraints: 1. SVG Elements: Use only the specified elements: rect, circle, ellipse, line, polyline, polygon, and short path (up to 5 commands). 2. Canvas Details: The SVG canvas is defined by a 512512 unit viewBox. Coordinates start at (0, 0) in the top-left and extend to (512, 512) at the bottom-right. 3. Element Stacking Order: The sequencing of SVG elements matters; elements defined later in the code will overlap earlier ones. 4. Colors: Use hexadecimal color values (e.g., #FF0000). For layers fully enclosed by others, differentiate with distinct colors. 5. Simplicity: Keep the SVG code simple and clear. 6. Realism: While using simple shapes, strive to create recognizable and proportionate representations of objects. Note: The SVG you create will serve as an initial draft using simple shapes rather than a fully polished final product with complex paths. Focus on creating a recognizable representation of the prompt using basic geometric forms. |
| Prompt Expansion |
| Expand the given text prompt to detail the abstract concept. Follow these steps in your response: 1. **Expand the Short Text Prompt**. Start by expanding the short text prompt into a more detailed description of the scene. Focus on talking about what objects appear in the scene, rather than merely talking the abstract idea. For example, for the short prompt “A spaceship flying in the sky”, expand it to “A silver spaceship soars through the galaxy, with distant planets and shimmering stars forming a vivid backdrop”. 2. **Object Breakdown and Component Analysis**. (a) For every object in the expanded prompt, add more details to describe the object. You can include color, size, shape, motion, status, or any other relevant details. For example, “A silver spaceship” can be expanded into “A silver spaceship with two large wings, ejecting flames from its thrusters”. (b) Then, break down each object into its individual components. For instance, the spaceship’s components could be “a body (rectangle), two triangular wings (polygon), a window (circle), and flames (polyline) emitting from the rear thrusters (rectangle)”. You need to list **ALL** parts of each object. If you ignore any part, the system will assume it’s not present in the scene. When listing components, explain how each component can be depicted using the specified SVG elements. 3. **Scene Layout and Composition**. Propose a logical and visually appealing layout for the final scene. For each object and each component, describe their positions on the canvas, relative/absolute sizes, colors, and relative spatial arrangement (e.g., the hand is connected to the arm, the moon is behind the mountain). |
| When expanding the prompt, follow these guidelines: 1. When expanding the short text prompt, avoid adding excessive new objects to the scene. Introduce additional objects only if they enhance the completeness of the scene. If the prompt mentions just a single object, you can choose to not introduce new objects and focus instead on enriching the description with more details about that object. 2. When add details to describe objects, the description can be detailed and vivid, but the language should be clear and concise. Avoid overly complex or ambiguous descriptions. 3. When breaking down objects into individual components, ensure you list all essential parts typically comprising that object, even if they are not explicitly mentioned in the initial object description. |
A Unicorn Example:
Input Text Prompt: “A unicorn is eating a carrot.”
Expanded Prompt: 1. Scene Description: “The pink unicorn is standing in side view. The unicorn’s mouth is open, bitting an orange carrot.” 2. Object Detail: • Unicorn: The unicorn, standing in side view, has a pink body and a yellow horn. Its brown tail swings, while its four legs stand still. It has a pink head and a small black eye. • Carrot: The carrot is orange with two green leaves at the top. 3. Component Breakdown: • Unicorn: horizontal pink body (ellipse), pink head (ellipse), thin neck (rectangle), four vertical legs (rectangles), brown curved tail (polyline), yello triangle horn atop the head (polygon), round eye (circle), mouth on the head (path) • Carrot: elongated orange triangle body (triangle), two small green triangular leaves (triangle) |
| Key Components Layout: 1. Unicorn (a) Body: An ellipse centered around (256, 256) with rx=90 and ry=60. The body is pink. (b) Head: An ellipse centered at (342, 166), 30 units wide and 25 units high, oriented to suggest the unicorn’s gaze forward. (c) Neck: A rectangle, 50 units long and 10 units wide, positioned at (312, 168) connecting the head and the body. (d) Legs: Four rectangles, each 10 units wide and 80 units tall, positioned at (185, 296), (220, 311), (293, 307), and (316, 296) to suggest stability. (e) Tail: A polyline starting from the back of the body ellipse at (168, 258) and curving to points (122, 298) and (142, 252) to suggest a flowing tail. (f) Horn: A polygon with points at (331, 140), (336, 110), and (341, 140) to represent the unicorn’s triangular horn. The horn is yellow. (g) Eye: A small circle with a radius of 5 units, placed at (352, 164) on the head. (h) Mouth: A small curved line positioned at (342, 178) on the head 2. Carrot (a) Body: An elongated triangle with points at (369, 180), (348, 200), and (356, 172). The carrot body is orange (#FFA500). (b) Leaves: Two small triangles positioned at the top of the carrot body, centered around (363, 174). The leaves are green (#00FF00). ##### Refer to the Unicorn example for response guidance and formatting. Avoiding any unnecessary dialogue in your response. Here is the text prompt: TEXT_PROMPT |
| SVG Script Generation |
| Write the SVG code following the expanded prompt and layout of key components, adhering to these rules: 1. SVG Elements: Use only the specified elements: rect, circle, ellipse, line, polyline, polygon, and short path (up to 5 commands). Other elements like text, Gradient, clipPath, etc., are not allowed. If there is path, the final command should be Z. If a path only contains a single command, you need to increase the stroke-width to make it visible. 2. Viewbox: The viewbox should be 512 by 512. 3. Stacking Order: Elements defined later will overlap earlier ones. So if there is a background, it should be defined first. 4. Colors: Use hexadecimal color values (e.g., #FF0000). For layers fully enclosed by others, differentiate with distinct colors. 5. Comments: Include concise phrase to explain the semantic meaning of each element. 6. Shape IDs: **Every** shape element should have a unique “id” starting with “path_num”. |
| The translation from the Unicorn’s expanded prompt to SVG code is provided below: ##### ⬇ <svg viewBox="0 0 512 512" xmlns="http://www.w3.org/2000/svg"> <!-- Body --> <ellipse id="path_1" cx="256" cy="256" rx="90" ry="60" fill="#ffc0cb"/> <!-- Legs --> <rect id="path_2" x="185" y="296" width="10" height="80" fill="#d3d3d3"/> <rect id="path_3" x="220" y="311" width="10" height="80" fill="#d3d3d3"/> <rect id="path_4" x="293" y="307" width="10" height="80" fill="#d3d3d3"/> <rect id="path_5" x="316" y="296" width="10" height="80" fill="#d3d3d3"/> <!-- Neck --> <rect id="path_6" x="312" y="168" width="10" height="50" fill="#ffc0cb"/> <!-- Head --> <ellipse id="path_7" cx="342" cy="166" rx="30" ry="25" fill="#ffc0cb"/> <!-- Eye --> <circle id="path_8" cx="352" cy="164" r="5" fill="#000000"/> <!-- Tail --> <polyline id="path_9" points="168,258 122,298 142,252" fill="none" stroke="#a52a2a" stroke-width="8"/> <!-- Horn --> <polygon id="path_10" points="331,140 336,110 341,140" fill="#ffff00"/> <!-- Unicorn mouth --> <path id="path_11" d="M 337 178 Q 342 183 347 178" fill="none" stroke="#000000" stroke-width="2"/> <!-- Carrot body --> <polygon id="path_12" points="369 180 348 200 356 172 369 180" fill="#ffa500"/> <!-- Carrot leaves --> <polygon id="path_13" points="363 174 364 163 373 168 363 174" fill="#00ff00"/> <polygon id="path_14" points="363 174 356 166 375 173 363 174" fill="#00ff00"/> </svg>< ##### |
| In your answer, avoid any unnecessary dialogue, and include the SVG code in the following format: ⬇ ‘‘‘svg svg_code ‘‘‘ |
| Visual Rectification |
| The SVG code you provide might have a critical issue: while it adheres to the text prompt, the rendered image could reveal real-world inconsistencies or visual oddities. For example: 1. Misalignments: The unicorn’s legs may appear detached from the body. 2. Hidden elements: The snowman’s arms could be hidden if they blend with the body due to identical colors and overlapping elements, making them indistinguishable. 3. Unrecognizable object: The SVG code includes a tiger, but the rendered image is unrecognizable due to a disorganized arrangement of shapes. 4. Disproportionate scaling: The squirrel’s tail might appear overly small compared to its body. 5. Color: If a shape is purely white and placed on a white background, it may seem invisible in the final image. If there is no background, try to avoid using white for the shape. 6. Incorrect path order: Incorrect path order can cause unintended overlaps, such as the face being completely covered by the hair or hat. |
These issues may not be evident in the SVG code but become apparent in the rendered image.
The provided image is rendered from your SVG code. You need to do the following: 1. First, carefully examine the image and SVG code to detect visual problems. Please list ALL the visual problems you find. If the image is severely flawed/unrecognizable, consider rewriting the entire SVG code. 2. Second, adjust the SVG code to correct these visual oddities, ensuring the final image appears more realistic and matches the expanded prompt. 3. When adding/deleting/modifying elements, ensure the IDs are unique and continuous, starting from “path_1”, “path_2”, etc. For example, if you delete “path_3”, rename “path_4” to “path_3” and “path_5” to “path_4” to maintain continuity. |
Your task is NOT to modify the SVG code to better match the image content, but to identify visual oddities in the image and suggest adjustments to the SVG code to correct them. You are not permitted to delete any SVG elements unless rewriting is involved.
In your answer, include the SVG code in the following format: ⬇ ‘‘‘svg svg_code ‘‘‘ |
| Editing |
| Analyze the given editing prompt to understand the user’s intention. Classify the editing instruction into one of (or a combination of) the following categories: 1. **Object Addition**: Adding a new object to the scene. 2. **Object Removal**: Removing an existing object from the scene. 3. **Object Modification**: Changing an existing object in the scene (e.g., color, size, position, pose, layout). |
| Follow these steps based on the type of editing instruction: • **Object Addition**: 1. Detailed Description: For each new object, add more details to describe the object. You can include color, size, shape, motion, status, or any other relevant details. For example, “A silver spaceship” can be expanded into “A silver spaceship with two large wings, ejecting flames from its thrusters”. 2. Component Breakdown: Break down each object into its individual components. For instance, the spaceship’s components could be “a body (rectangle), two triangular wings (polygon), a window (circle), and flames (polyline) emitting from the rear thrusters (rectangle)”. You need to list **ALL** parts of each object. If you ignore any part, the system will assume it’s not present in the scene. When listing components, explain how each component can be depicted using the specified SVG elements. 3. Global Layout: Propose a global layout for each new object, i.e., describing its spatial relationship to existing elements. 4. Local Components Layout: Propose a local components layout, describing arrangement of its components, including their relative sizes and positions. 5. Stacking Order: Specify the layering order of new elements, especially for overlapping objects, to ensure the correct visual effect. • **Object Removal**: 1. Identify the object(s) to be removed with precise descriptions. 2. Specify any adjustments needed for remaining elements to maintain scene coherence. • **Object Modification**: 1. Identify the specific part(s) of the object that need to be changed. 2. Describe modifications in detail, including exact size, colors (in hexadecimal), and positions where applicable. 3. For complex modifications, consider treating them as a combination of object removal and addition. |
| Guidelines for expanding the prompt: 1. When add details to describe objects, the description can be detailed and vivid, but the language should be clear and concise. Avoid overly complex or ambiguous descriptions. 2. When breaking down objects into individual components, ensure you list all essential parts typically comprising that object, even if they are not explicitly mentioned in the initial object description. 3. For object modifications, provide exact specifications (e.g., “Increase the unicorn’s horn length by 20 units”). 4. Consider the overall composition and balance of the scene when adding or modifying elements. |
| Operation Summary: In the original SVG code, each shape has an id attribute. After generating the edited SVG code, summarize the operations performed in the following format: 1. Element Modification: [id_1, id_2, …] (for elements that were modified but kept the same stacking order) 2. Element Removal: [id_1, id_2, …] (for elements that were removed) 3. Element Addition: start_path_id, [id_1, id_2, …] (for newly added elements; start_path_id is the id of the path in the original SVG code, indicating that new elements should be inserted after the element with id start_path_id. If inserting at the beginning, set start_path_id to an empty string “”.) Note: The ids in the lists refer to the ids from the original SVG code, not the edited SVG code. |